Indium has 2 isotopes. The 113-In isotope has a mass of 112.90406184 amu and a percent abundance of 4.290%. What is the mass (amu) of the other isotope?
Answer #1

Given that,

Mass of first isotope 113 In is 112.90406184 amu

Its percentage abundance = 4.290 % = 0.04290

Mass of second 115 In isotope = x (say)

Its abundance = 100 - 0.04290 = 0.9571

Atomic mass of indium = 114.818 amu

But atomic mass = (Mass of first isotope *its abundance)+ (Mass of 2nd isotope *its abundance)

114.818 amu = (112.90406184 *0.04290 ) + ( x * 0.9571)

x = 114.9 amu

Thus, the mass of 2nd isotope = 114.9 amu